OpenStudy (anonymous):

identify a pattern and find the next number in the pattern. -48,-24,-12,-6?

OpenStudy (anonymous):

pattern is multiply by two next # (-48)(2)=-96 then#(-96)(2)=-192 patter is divide by two going the other way next#(-6)/2=-3 next#(-3)/2=-3/2

OpenStudy (anonymous):

thank youuuu i have another equation exactly the same can you help?

OpenStudy (anonymous):

-0.6 , -3 , -15 , -75?

OpenStudy (anonymous):

multiply by 5

OpenStudy (anonymous):

(-75)(5)=-375 (-375)(5)=-1875